# Bayesian Perspective for Orientation Determination in Cryo-EM with Application to Structural Heterogeneity Analysis

### v2

@@ Appendix A Full cryo-EM model
Appendix B Alignment between two noisy volumes
@@
From (B.1), the unknown volume $V$ can be expressed as:
$V=g_{1}^{-1}\circ(y_{1}-\epsilon_{1}).$ (B.2)
Substituting (B.2) into (B.1), yields:
$y_{2}$ $=g_{2}\circ V+\epsilon_{2}$ (B.3)
$=g_{2}\circ\left(g_{1}^{-1}\circ(y_{1}-\epsilon_{1})\right)+\epsilon_{2}$ (B.4)
$=(g_{2}\cdot g_{1}^{-1})\circ y_{1}-(g_{2}\cdot g_{1}^{-1})\circ\epsilon_{1}+\epsilon_{2}$ (B.5)
$=(g_{2}\cdot g_{1}^{-1})\circ y_{1}+\epsilon,$ (B.6)
where (B.5) follows from the associativity of group actions, and in(B.6), we have used $\epsilon\triangleq(g_{2}\cdot g_{1}^{-1})\circ\epsilon_{1}+\epsilon_{2}$. Since the noise statistics are invariant under group actions, it follows that $\epsilon\sim\mathcal{N}(0,2\Sigma_{d\times d})$.
@@
Appendix C Proof of Proposition 2.1
@@
$\hat{g}_{\mathsf{MMSE}}=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,min}}\int_{\mathsf{SO(3)}}\left\|g-\hat{g}\right\|_{F}^{2}{\mathbb{P}}_{\Lambda}(g|y,{V})\mathrm{d}g.$ (C.1)
@@
$\hat{g}_{\mathsf{MMSE}}$ $=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\int_{\mathsf{SO(3)}}\langle\hat{g},g\rangle{\mathbb{P}}_{\Lambda}(g|y,{V})\mathrm{d}g$ (C.2)
$=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\langle\hat{g},\hat{g}_{\mathsf{relax}}\rangle.$ (C.3)
As a consequence, (C.3) is equal to the rotation $\hat{g}\in\mathsf{SO}(3)$ that has the highest correlation with the relexed estimator $\hat{g}_{\mathsf{relax}}$ as in (2.12). It follows that
$\hat{g}_{\mathsf{MMSE}}=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\langle\hat{g},\hat{g}_{\mathsf{relax}}\rangle=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,min}}\left\|\hat{g}-\hat{g}_{\mathsf{relax}}\right\|_{F}^{2}$ (C.4)
@@
Appendix D Proof of Proposition 2.2
@@
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MAP}}=\hat{g}_{\mathsf{MLE}}.$ (D.1)
@@
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MMSE}}=\hat{g}_{\mathsf{MLE}}.$ (D.2)
For (D.1), the argument is straightforward. Since the prior density $\Lambda(g)$ is bounded below by some positive constant, the term $\log(\mathrm{d}\Lambda(g))$ becomes negligible as $\sigma\to 0$ compared to the other term. Consequently, the MAP estimator converges to the MLE in this limit.
For (D.2), we apply the following theorem [42, Corollary 5.11]:
Theorem D.1.
@@
$\theta^{*}=\underset{\theta\in\Theta}{\operatorname*{arg\,max}}\ h(\theta),$ (D.3)
@@
$\underset{\lambda\to\infty}{\lim}\frac{\int_{\Theta}\theta e^{\lambda h(\theta)}\pi(\theta)\mathrm{d}\theta}{\int_{\Theta}e^{\lambda h(\theta)}\pi(\theta)\mathrm{d}\theta}=\theta^{*},$ (D.4)
@@
$=\underset{g\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\ h(g)=\hat{g}_{\mathsf{MLE}}.$ (D.5)
@@
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MMSE}}=\hat{g}_{\mathsf{MLE}}.$ (D.6)
Appendix E Isotropic Gaussian distribution of SO(3) rotations
@@
$f(\omega)=\frac{1-\cos\omega}{\pi}\sum_{\ell=0}^{\infty}(2\ell+1)e^{-\ell(\ell+1)\eta^{2}}\frac{\sin\left((\ell+1/2)\omega\right)}{\sin(\omega/2)}.$ (E.1)
Notably, the uniform distribution on $\mathsf{SO}(3)$, denoted $\mathcal{U}_{\mathsf{SO}(3)}$, corresponds to uniformly sampled axes and is characterized by the density function $f(\omega)=\frac{1-\cos\omega}{\pi}$, which represents the limiting case of $\eta\to\infty$ in (E.1).
@@
$f(\omega)\simeq\frac{1-\cos\omega}{\sqrt{\pi}}\eta^{-\frac{3}{2}}e^{\left(\frac{\eta}{4}-\frac{\left(\omega/2\right)^{2}}{\eta}\right)}\left(\frac{\omega-e^{-\pi^{2}/\eta}\left(\left(\omega-2\pi\right)e^{\pi\omega/\eta}+\left(\omega+2\pi\right)e^{-\pi\omega/\eta}\right)}{2\sin\left(\frac{\omega}{2}\right)}\right).$ (E.2)
@@
Appendix F Proof of Proposition 4.1
@@
$\hat{V}^{(t+1)}=\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\|y_{i}-\left(g\circ V\right)\right\|^{2}\mathrm{d}g,$ (F.1)
@@
$p_{i}^{(t)}(g)=\frac{\exp\Big(-\left\|y_{i}-\left(g\circ\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big)}{\int_{g\in\mathsf{SO}(3)}\exp\Big(-\left\|y_{i}-\left(g\circ\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big)\mathrm{d}g}.$ (F.2)
@@
$\left\|y_{i}-\left(g\circ V\right)\right\|^{2}=\left\|g^{-1}\circ{y_{i}}-V\right\|^{2}=\left\|g^{-1}\circ y_{i}\right\|^{2}+\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}.$ (F.3)
Thus, substituting (F.3) into (F.1), leads to:
$\hat{V}^{(t+1)}=\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|g^{-1}\circ y_{i}\right\|^{2}+\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}\right\}\mathrm{d}g.$ (F.4)
As the term $\left\|g\circ y_{i}\right\|^{2}$ is independent of $V$, (F.4) can be simplified to:
$\hat{V}^{(t+1)}=\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}\right\}\mathrm{d}g.$ (F.5)
Let us denote by $h(V)$ the right hand side of (F.5):
$h(V)\triangleq\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}\right\}\mathrm{d}g.$ (F.6)
@@
$\frac{\partial}{\partial V}h(V)=2\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left(V-g^{-1}\circ y_{i}\right)\mathrm{d}g=0.$ (F.7)
As $\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)=1$ for every $i$, we can simplify (F.7) to obtain:
@@
$=$ $\frac{1}{M}\sum_{i=0}^{M-1}\big(\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\cdot g^{-1}\mathrm{d}g\big)\circ y_{i}.$ (F.8)
@@
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}\circ y_{i},$ (F.9)

### v3

@@ Appendix A Full cryo-EM model
Appendix B Bayes pose estimation with rotations and shifts
To account for in-plane shifts, we model the latent pose as $\phi=(g,t)\in\mathsf{SO}(3)\times\mathbb{R}^{2}$, where $g$ denotes the 3D rotation and $t$ denotes the in-plane translation. Given a volume $V$, the forward model can be written as
$p\!\left(y\mid g,t\right)\propto\exp\!\Big(-\tfrac{1}{2}\,\|y-\mathcal{A}(g,t)V\|_{\Sigma^{-1}}^{2}\Big),$ (B.1)
where $\mathcal{A}(g,t)$ is the cryo-EM/ET forward operator including both rotation and shift (e.g., in the Fourier domain the shift acts via the phase factor $e^{-2\pi i\,\omega\cdot t}$; cf. Appendix A). We assume priors $\Lambda(g)$ on rotations and $\pi(t)$ on shifts. The joint posterior then takes the form
$p(g,t\mid y)\propto p(y\mid g,t)\,\Lambda(g)\,\pi(t).$ (B.2)
Additive loss function.
To define a Bayes estimator, we specify a loss function on $\mathsf{SO}(3)\times\mathbb{R}^{2}$. Throughout, we adopt an additive pose loss,
[…]
which penalizes rotation and translation errors separately (optionally with application-dependent weights to reconcile units). Under (B.3), the Bayes pose estimator is defined by
$(\widehat{g}_{\mathrm{Bayes}},\widehat{t}_{\mathrm{Bayes}})\in\operatorname*{arg\,min}_{\widehat{g}\in\mathsf{SO}(3),\,\widehat{t}\in\mathbb{R}^{2}}\ \mathbb{E}\!\left[\mathcal{L}\big((g,t),(\widehat{g},\widehat{t}\,)\big)\,\big|\,y\right],$ (B.4)
where the expectation is taken with respect to the joint posterior (B.2). Because the loss is additive, the minimization in (B.4) decomposes into two marginal Bayes estimators:
$\widehat{g}_{\mathrm{Bayes}}$ $=\operatorname*{arg\,min}_{\widehat{g}\in\mathsf{SO}(3)}\ \mathbb{E}\!\left[\mathcal{L}_{g}(g,\widehat{g})\mid y\right],$ (B.5)
$\widehat{t}_{\mathrm{Bayes}}$ $=\operatorname*{arg\,min}_{\widehat{t}\in\mathbb{R}^{2}}\ \mathbb{E}\!\left[\mathcal{L}_{t}(t,\widehat{t}\,)\mid y\right].$ (B.6)
Closed-form MMSE rules under squared loss.
For translations, choosing Euclidean squared loss $\mathcal{L}_{t}(t,\widehat{t}\,)=\|t-\widehat{t}\,\|_{2}^{2}$ yields the posterior mean (similarly to (2.13))
$\widehat{t}_{\mathrm{MMSE}}=\mathbb{E}[t\mid y]=\int_{\mathsf{SO}(3)}\int_{\mathbb{R}^{2}}t\,p(g,t\mid y)\,\mathrm{d}t\,\mathrm{d}g,$ (B.7)
where $p(g,t\mid y)$ is defined through (B.2). For rotations, when $g$ is represented by its $3\times 3$ matrix and we take the squared chordal/Frobenius loss $\mathcal{L}_{g}(g,\widehat{g})=\|g-\widehat{g}\|_{F}^{2}$, the corresponding Bayes estimator is obtained by first forming the relaxed posterior mean matrix (similarly to (2.13))
$\widehat{g}_{\mathrm{relax}}=\mathbb{E}[g\mid y]=\int_{\mathsf{SO}(3)}\int_{\mathbb{R}^{2}}g\,p(g,t\mid y)\,\mathrm{d}t\,\mathrm{d}g,$ (B.8)
and then projecting it onto $\mathsf{SO}(3)$ via the Orthogonal Procrustes map (similarly to (2.17)),
$\widehat{g}_{\mathrm{MMSE}}=\operatorname*{arg\,min}_{\Omega\in\mathsf{SO}(3)}\ \big\|\widehat{g}_{\mathrm{relax}}-\Omega\big\|_{F}.$ (B.9)
which returns the closest valid rotation matrix (in Frobenius norm) to $\widehat{g}_{\mathrm{relax}}$.
Appendix C Alignment between two noisy volumes
@@
From (C.1), the unknown volume $V$ can be expressed as:
$V=g_{1}^{-1}\circ(y_{1}-\epsilon_{1}).$ (C.2)
Substituting (C.2) into (C.1), yields:
$y_{2}$ $=g_{2}\circ V+\epsilon_{2}$ (C.3)
$=g_{2}\circ\left(g_{1}^{-1}\circ(y_{1}-\epsilon_{1})\right)+\epsilon_{2}$ (C.4)
$=(g_{2}\cdot g_{1}^{-1})\circ y_{1}-(g_{2}\cdot g_{1}^{-1})\circ\epsilon_{1}+\epsilon_{2}$ (C.5)
$=(g_{2}\cdot g_{1}^{-1})\circ y_{1}+\epsilon,$ (C.6)
where (C.5) follows from the associativity of group actions, and in(C.6), we have used $\epsilon\triangleq(g_{2}\cdot g_{1}^{-1})\circ\epsilon_{1}+\epsilon_{2}$. Since the noise statistics are invariant under group actions, it follows that $\epsilon\sim\mathcal{N}(0,2\Sigma_{d\times d})$.
@@
Appendix D Proof of Proposition 2.1
@@
$\hat{g}_{\mathsf{MMSE}}=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,min}}\int_{\mathsf{SO(3)}}\left\|g-\hat{g}\right\|_{F}^{2}{\mathbb{P}}_{\Lambda}(g|y,{V})\mathrm{d}g.$ (D.1)
@@
$\hat{g}_{\mathsf{MMSE}}$ $=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\int_{\mathsf{SO(3)}}\langle\hat{g},g\rangle{\mathbb{P}}_{\Lambda}(g|y,{V})\mathrm{d}g$ (D.2)
$=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\langle\hat{g},\hat{g}_{\mathsf{relax}}\rangle.$ (D.3)
As a consequence, (D.3) is equal to the rotation $\hat{g}\in\mathsf{SO}(3)$ that has the highest correlation with the relexed estimator $\hat{g}_{\mathsf{relax}}$ as in (2.13). It follows that
$\hat{g}_{\mathsf{MMSE}}=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\langle\hat{g},\hat{g}_{\mathsf{relax}}\rangle=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,min}}\left\|\hat{g}-\hat{g}_{\mathsf{relax}}\right\|_{F}^{2}$ (D.4)
@@
Appendix E Proof of Proposition 2.2
@@
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MAP}}=\hat{g}_{\mathsf{MLE}}.$ (E.1)
@@
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MMSE}}=\hat{g}_{\mathsf{MLE}}.$ (E.2)
For (E.1), the argument is straightforward. Since the prior density $\Lambda(g)$ is bounded below by some positive constant, the term $\log(\mathrm{d}\Lambda(g))$ becomes negligible as $\sigma\to 0$ compared to the other term. Consequently, the MAP estimator converges to the MLE in this limit.
For (E.2), we apply the following theorem [45, Corollary 5.11]:
Theorem E.1.
@@
$\theta^{*}=\underset{\theta\in\Theta}{\operatorname*{arg\,max}}\ h(\theta),$ (E.3)
@@
$\underset{\lambda\to\infty}{\lim}\frac{\int_{\Theta}\theta e^{\lambda h(\theta)}\pi(\theta)\mathrm{d}\theta}{\int_{\Theta}e^{\lambda h(\theta)}\pi(\theta)\mathrm{d}\theta}=\theta^{*},$ (E.4)
@@
$=\underset{g\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\ h(g)=\hat{g}_{\mathsf{MLE}}.$ (E.5)
@@
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MMSE}}=\hat{g}_{\mathsf{MLE}}.$ (E.6)
Appendix F Non-uniform distribution of rotations
F.1 Isotropic Gaussian distribution of SO(3) rotations
@@
$f(\omega)=\frac{1-\cos\omega}{\pi}\sum_{\ell=0}^{\infty}(2\ell+1)e^{-\ell(\ell+1)\eta^{2}}\frac{\sin\left((\ell+1/2)\omega\right)}{\sin(\omega/2)}.$ (F.1)
Notably, the uniform distribution on $\mathsf{SO}(3)$, denoted $\mathcal{U}_{\mathsf{SO}(3)}$, corresponds to uniformly sampled axes and is characterized by the density function $f(\omega)=\frac{1-\cos\omega}{\pi}$, which represents the limiting case of $\eta\to\infty$ in (F.1).
To provide intuition for the spread induced by $\eta$, consider the concentrated regime $\eta\ll 1$. In this regime, one may use the standard small-angle approximation in which a rotation is represented by a rotation vector $r\in\mathbb{R}^{3}$ with
$r\sim\mathcal{N}(0,\eta^{2}I_{3}),\qquad\omega\approx\|r\|,$ (F.2)
so that $\omega$ is approximately Maxwell distributed with scale parameter $\eta$ [25]. Consequently, its standard deviation is
$\mathrm{Std}(\omega)\approx\eta\,\sqrt{3-\frac{8}{\pi}}=\eta\,\sqrt{\frac{3\pi-8}{\pi}}\approx 0.67\,\eta\quad\text{(radians)}.$ (F.3)
For example, $\eta=0.1$ corresponds to $\mathrm{Std}(\omega)\approx 0.067$ rad $\approx 3.8^{\circ}$.
@@
$f(\omega)\simeq\frac{1-\cos\omega}{\sqrt{\pi}}\eta^{-\frac{3}{2}}e^{\left(\frac{\eta}{4}-\frac{\left(\omega/2\right)^{2}}{\eta}\right)}\left(\frac{\omega-e^{-\pi^{2}/\eta}\left(\left(\omega-2\pi\right)e^{\pi\omega/\eta}+\left(\omega+2\pi\right)e^{-\pi\omega/\eta}\right)}{2\sin\left(\frac{\omega}{2}\right)}\right).$ (F.4)
@@
F.2 Estimating the rotational prior
A key practical ingredient in Bayesian pose estimation is the prior over rotations. From an algorithmic standpoint, two complementary approaches are currently being developed to handle unknown viewing-angle distributions. First, building on EM-type refinement, recent work proposes to introduce an explicit parameterization of the orientation distribution over $\mathsf{SO}(3)$ and to jointly optimize the structure and the viewing-angle distribution in the M-step, effectively recovering the rotational prior from the data [62]. This perspective generalizes Bayesian EM refinement pipelines such as RELION by treating the pose prior as an estimable object that can be learned jointly with the structure during refinement [49]. Second, method-of-moments approaches relax the uniformity assumption and aim to recover both the structure and the unknown viewing-angle distribution from low-order statistics, typically via alternating optimization between the two components [50, 28].
In practice, we recommend starting with a uniform (or weak) prior when little is known about the rotational prior, and then optionally estimating a smoothed viewing-angle distribution from an initial refinement and plugging it into the MMSE estimator; simple sensitivity checks (uniform vs. learned priors, and varying regularization) can help diagnose the effect of preferred orientations. We emphasize, however, that principled estimation of the rotational prior, and a full theory of robustness to prior misspecification in cryo-EM/ET, remains an important open direction for future work.
Appendix G Proof of Proposition 4.1
@@ F.2 Estimating the rotational prior
$\hat{V}^{(t+1)}=\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\|y_{i}-\left(g\circ V\right)\right\|^{2}\mathrm{d}g,$ (G.1)
@@
$p_{i}^{(t)}(g)=\frac{\exp\Big(-\left\|y_{i}-\left(g\circ\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big)}{\int_{g\in\mathsf{SO}(3)}\exp\Big(-\left\|y_{i}-\left(g\circ\hat{V}^{(t)}\right)\right\|^{2}/2\sigma^{2}\Big)\mathrm{d}g}.$ (G.2)
@@
$\left\|y_{i}-\left(g\circ V\right)\right\|^{2}=\left\|g^{-1}\circ{y_{i}}-V\right\|^{2}=\left\|g^{-1}\circ y_{i}\right\|^{2}+\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}.$ (G.3)
Thus, substituting (G.3) into (G.1), leads to:
$\hat{V}^{(t+1)}=\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|g^{-1}\circ y_{i}\right\|^{2}+\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}\right\}\mathrm{d}g.$ (G.4)
As the term $\left\|g\circ y_{i}\right\|^{2}$ is independent of $V$, (G.4) can be simplified to:
$\hat{V}^{(t+1)}=\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}\right\}\mathrm{d}g.$ (G.5)
Let us denote by $h(V)$ the right hand side of (G.5):
$h(V)\triangleq\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\{\left\|V\right\|^{2}-2{\,\langle{g^{-1}\circ y_{i}},V\rangle}\right\}\mathrm{d}g.$ (G.6)
@@
$\frac{\partial}{\partial V}h(V)=2\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left(V-g^{-1}\circ y_{i}\right)\mathrm{d}g=0.$ (G.7)
As $\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)=1$ for every $i$, we can simplify (G.7) to obtain:
@@
$=$ $\frac{1}{M}\sum_{i=0}^{M-1}\big(\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\cdot g^{-1}\mathrm{d}g\big)\circ y_{i}.$ (G.8)
@@
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}\circ y_{i},$ (G.9)